OpenStudy (anonymous):

Find the probability of picking a blue marble and a yellow marble when 2 marbles are picked (without replacement)from a bag containing 4 blue and 5 yellow marbles.

OpenStudy (anonymous):

Write a proportion

OpenStudy (anonymous):

how am i suppose to do that?

OpenStudy (anonymous):

Do you know how to write a proportion? It's two fractions equal to each other. If you need more help, go on Kahn Academy

jimthompson5910 (jim_thompson5910):

4 blue + 5 yellow = 9 total the probability of picking a blue marble on the first draw is (# of blue)/(# total) = 4/9

jimthompson5910 (jim_thompson5910):

no replacements are made, so after the first draw we have 9-1 = 8 marbles left over. There are still 5 yellow marbles the probability of picking yellow on the second draw is (# of yellow)/(# of marbles left) = 5/8

OpenStudy (anonymous):

what about the yellow

jimthompson5910 (jim_thompson5910):

multiply the two fractions to get the answer

OpenStudy (anonymous):

oh